10. A piggy bank contains hundred 50p coins, fifty ₹1 coins, twenty ₹2 coins and ten ₹5 coins. If it is equally likely that one of the coins will fall out when the bank is turned upside down, what is the probability that the coin

(i) will be a 50 p coin?

(ii) will not be a ₹5 coin?

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
Total no. of coins = 100+50+20+10 = 180 P(E) = (Number of favourable outcomes/ Total number of outcomes) (i) Total number of 50 p coin = 100 P (50 p coin) = 100/180 = 5/9 = 0.55 (ii) Total number of ₹5 coin = 10 P (₹5 coin) = 10/180 = 1/18 = 0.055 ∴ P (not ₹5 coin) = 1-P (₹5 coin) = 1-0.055 = 0.945